OpenStudy (anonymous):

(x+5/x-5) / (x^2-25/x-5)

OpenStudy (watchmath):

\(\frac{x+5}{x-5}\cdot\frac{x-5}{(x-5)(x+5)}=\frac{1}{x-5}\)